OpenStudy (lindsey11x16):

how does sec^4x - tan^4x go with sec^2x + tan^2x

OpenStudy (mayankdevnani):

sec^4x-tan^4x =(sec^2x+tan^2x)(sec^2x-tan^2x) = sec^2 x+ tan ^2 x as sec^2x -tan^2 x = 1

OpenStudy (lindsey11x16):

?????????????????

OpenStudy (mayankdevnani):

sec⁴x - tan⁴x = sec²x + tan²x → 1/(cos⁴x) - (sin⁴x)/(cos⁴x) = 1/(cos²x) + (sin²x)/(cos²x) let us now multiply both sides by cos⁴x : (cos⁴x)/(cos⁴x) - (sin⁴x)(cos⁴x) /(cos⁴x) = (cos⁴x) /(cos²x) + (sin²x)(cos⁴x) /(cos²x) → 1 - sin⁴x = cos²x + sin²x cos²x and let now express everything as a function of sinx: 1 - sin⁴x = 1 - sin²x + sin²x (1 - sin²x) → 1 - sin⁴x = 1 - sin²x + sin²x - sin⁴x → 1 - sin⁴x = 1 - sin⁴x (verified)

OpenStudy (lindsey11x16):

how did the negative in sec⁴x - tan⁴x become positive in sec²x + tan²x